Butyric acid contains C, H and O. A 4.24 mg sample of it is burned which gives 8.45 mg CO2 and 3.46 mg H2O. What is the molecular formula of Butyric acid if it`s molecular mass is 88 amu?

Grade:11

1 Answers

Vikas TU
7 years ago
number of moles = mass / molar mass
molar mass of CO2 = 44 g/ mole
0.00845 g of CO2 has 0.008 /44 = 0.00019 moles of CO2
there is 1 mole of C in CO2 and all the C from the compound becomes CO2 
so moles of C in the compound = 0.00019 moles
mass of C = 0.00019 x 12 = 0.00231 g

molar mass of H2O = 18 g/ mole
0.00346 g of H2O has 0.00346 / 18 = 0.00019 moles of H2O
there are 2 moles of H in H2O so moles of H in the compound = 0.00038 moles
mass of H = 0.00038 x 1.0079 = 0.0004 g

mass of H + C = 0.00269 g
mass of sample = 0.00424 g
mass of O by difference = 0.00155 g
moles of O = 0.00155 /16 = 0.00010 moles 

molar ratio of C : H : O = 0.00019 : 0.00038 : 0.00010 
smallest number 0.00010
divide the ratio by the smallest number we get 
molar ratio of C : H : O = 2.0 : 4.0 : 1.0 
empirical formula is C2H4O